OpenStudy (anonymous):

if a^2 - b^2 = (a + b)(a - b) does that mean a^2 + b^2 = (a - b)(a + b) ?? Thanks :)

OpenStudy (anonymous):

no (a-b)(a+b) is this a^2-b^2

OpenStudy (anonymous):

\[x*y=y*x\]So (a+b)(a-b) = (a-b)(a+b)

OpenStudy (anonymous):

So it's the commutative law of multiplication?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

thanks man :)